OpenStudy (anonymous):

Find te perimeter of a parallelogram. Give and example of this and include the formula you used.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@anonymous_user @phi

OpenStudy (anonymous):

Please help!! Dx

OpenStudy (anonymous_user):

Do you have a picture?

OpenStudy (anonymous):

No I have to answer it just like that.

OpenStudy (anonymous):

I mean you can make it up. Like just give an example and stuff.

OpenStudy (anonymous_user):

@SuperRbow15

OpenStudy (aum):



OpenStudy (aum):

Parallelograms have opposite sides equal. Perimeter = a + b + a + b = 2a + 2b = 2(a + b) You can pick two numbers. Say a = 7, b = 4 Then perimeter = 2(7 + 4) = 2(11) = 22

OpenStudy (anonymous):

Can you like make it simpler? I need to like say this to my teacher on the phone. And um she might not understand.

OpenStudy (phi):

Does this part make sense: opposite sides are equal

OpenStudy (phi):

for example a "box" i.e. a rectangle 

OpenStudy (phi):

the 2 short sides are the same length the 2 long sides (top and bottom) are the same length that is true for all parallelograms

OpenStudy (phi):

the "perimeter" is the distance if you walked along the lines, all the way round

OpenStudy (anonymous):

No I don't mean that. The old example was perfect but too many numbers.

OpenStudy (anonymous):

Actually nvm xD I get it.

OpenStudy (anonymous):

Can you help me with a few more questions? Like 2 more?

OpenStudy (phi):

if you make a new post

OpenStudy (anonymous):

Oh yeah i'll do that.